OpenStudy (anonymous):

Exact ODE: sin(x)cos(y) dx + cos(x)sin(y) dy = 0

OpenStudy (loser66):

- sinx cos y dx = cos x sin y dy \[-\frac{sinx}{cosx}dx =\frac{siny}{cosy}dy\]so, y = -x +C